OpenStudy (anonymous):

Find the derivatives of the function F(x) = integral from ((-2x+1) to (2x+1) of (t^3)/(sqrt(t^2+9)))dt

myininaya (myininaya):

\[\frac{d}{dt}(\int\limits_{-2x+1}^{2x+1}\frac{t^3}{\sqrt{t^2+9}} dt)?\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

have fun break it up into two pieces and then use the chain rule twice. in other words start with \[\int_{-2x+1}^a\frac{t^3}{\sqrt{t^2+9}}dt+\int_a^{2x+1}\frac{t^3}{\sqrt{t^2+9}}\] then \[-\int_a^{-2x+1}\frac{t^3}{\sqrt{t^2+9}}dt+\int_a^{2x+1}\frac{t^3}{\sqrt{t^2+9}}\]

myininaya (myininaya):

\[\text{ Let } g(t)=\frac{t^3}{\sqrt{t^2+9}} \text{ and let G be antiderivative of g}\] \[\frac{d}{dt}(G(t)|_{-2x+1}^{2x+1})=\frac{d}{dt}(G(2x+1)-G(-2x+1))\] \[(2x+1)'g(2x+1)-(-2x+1)'g(-2x+1)\]

myininaya (myininaya):

\[2g(2x+1)+2g(-2x+1)\]

OpenStudy (anonymous):

for the first one replace each t by \[-2x+1\] and multiply the result by -2 to get \[\frac{2(-2x+1)^3}{\sqrt{(-2x+1)^2+9}}\]

myininaya (myininaya):

\[2[\frac{(2x+1)^3}{\sqrt{(2x+1)^2+9}}+\frac{(-2x+1)^3}{\sqrt{(-2x+1)^2+9}}]\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

nvm sorry i thought i entered it wrong

OpenStudy (anonymous):

we are using two different methods but it amounts to the same thing

myininaya (myininaya):

goodnight i'm not simplifying that mess

OpenStudy (anonymous):

haha ok i got that far

OpenStudy (anonymous):

good night irene

myininaya (myininaya):

goodnight jack

OpenStudy (anonymous):

any tips for simplification?

OpenStudy (anonymous):

no simplification. leave it